The sum of two numbers is 15. Three times one of the numbers is 11 less than five times the other. Find the numbers.
Answer by iryneigh(1)

let x=1st number
let y= 2nd number
x + y = 15
x = 15 - y eq. 1
3x = 5y - 11 eq. 2
Substitute eq. 1 to eq. 2
3(15 - y) = 5y -11
45 - 3y = 5y -11
45 + 11 = 5y + 3y
56 = 8y
y = 56/8
y = 7
Subst. to eq. 1
x = 15 - 7
x = 8